Sarah drove her boat at 5 miles per hour for 45 minutes and at 4 miles per hour for 30 minutes. What was the total distance, in miles, that Sarah drove her boat?
A)2 miles
B)5 3/4 miles
C)9 miles
D)16 1/2 miles
I don't understand how to do this? Can you help?

Answer by Alan3354(69443)

Sarah drove her boat at 5 miles per hour for 45 minutes and at 4 miles per hour for 30 minutes. What was the total distance, in miles, that Sarah drove her boat?
------------------
change minutes to hours
------------------
Sarah drove her boat at 5 miles per hour for 3/4 hour and at 4 miles per hour for 1/2 hour. What was the total distance, in miles, that Sarah drove her boat?
d = r*t
d = 5*(3/4) + 4*(1/2)
d = 15/4 + 2 = 23/4 miles
d = 5 3/4 miles